A body at rest is given an initial uniform acceleration of 8.0m/s square for the next 20s. The body maintains the speed attained for 60s after which it is brought to rest in 20s. Draw the velocity-time graph of the motion using the information given above. Using the graph, calculate the (I) maximum speed attained during the motion. (ii) average retardation as the body is being brought to rest. (iii) total distance travelled during the first 50s. (iv) average speed during the same interval as in (ii).
1
Expert's answer
2019-08-27T09:35:21-0400

The velocity-time graph of the motion 




i) The maximum speed attained during the motion:


"V=160\frac{m}{s}"

ii) The average retardation as the body is being brought to rest:


"a=8\frac{m}{s^2}"

iii) The total distance traveled during the first 50 s:


"s=0.5(160)(20)+(160)(50-20)=6400\ m"

iv) The average speed during the same interval as in (ii):


"v_a=\frac{s}{t}=\frac{6400}{50}=128\frac{m}{s}"